Anew car worth 18,000 is depreciating by value by $3,000 per year. X should represent the # of years since the cars purchase, write an expression for the cars value using the given information.
original value depreciates by (/ - x +)
by 3,000 per year for x years

Answer by checkley77(12844)

18,000-(3,000x)
For 5 years:
18,000-3,000*5=18,000-15,000=3,000 is the car value after 5 years.